OpenStudy (anonymous):

Use the balanced chemical equation to fill in the blanks. All answers should include only 2 numbers. 3Ag2*S+2Al -> 6Ag +Al2*S3* *all numbers with star means that they are small and they just cant be subscripted thnks* Question: Safal used 7.8 moles of aluminum and _____moles of silver sulfide to complete his reaction. When he was finished he had created _____moles of silver and _____moles of aluminum sulfide. PLEASE HELP!!MY HEAD IS BURSTING

OpenStudy (anonymous):

We have the following chemical equation\[\rm 3 Ag_2 S + 2 Al \rightarrow 6 Ag + Al_2S_3\] From this equation, we know all the mole relations between each species involved. If we use 7.8 moles of aluminum (Al), we need to determine the ratio of the actual number of moles used in the laboratory to the number of moles in the chemical reaction. \[r = {7.8 \over 2}\] Now, we just multiply through our chemical equation with r to find the corresponding number of moles of the other species, when 7.8 moles of Aluminium are used. \[\rm (3 \cdot r) Ag_2 S + (2 \cdot r) Al \rightarrow (6 \cdot r) Ag + (r) Al_2S_3\] This is because if 2 moles of Aluminum, when reacted with 3 moles of Silver Sulfide, yields 6 moles of Silver and 1 moles of Aluminum Sulfide. Now, say we used 8 moles of Aluminum in the reaction. We would need 4 times as much Silver Sulfide (12 moles), and the reaction would yield 4 times as much Silver (24 moles) and 4 times as much Aluminum Sulfide (4 moles). The ratio, r, in this case is 4 because 8/2 = 4.